A body moving at 12 m/s accelerates to a velocity of 18 m/s</span>. If it covers a distance of 331 m during this time, what is the acceleration?  


1
Expert's answer
2021-03-16T11:37:44-0400
"a=\dfrac{v_f^2-v_i^2}{2d},""a=\dfrac{(18\ \dfrac{m}{s})^2-(12\ \dfrac{m}{s})^2}{2\cdot331\ m}=0.27\ \dfrac{m}{s^2}."